OpenStudy (mony01):

Before reaching cruising speed, when the car was speeding up from 0 to 50 mph, was the forward force greater than, equal to, or less than the backwards force? Explain

OpenStudy (anonymous):

Isn't this Physics?

OpenStudy (mony01):

yea are you able to help me? @Jesstho.-.

OpenStudy (swissgirl):

acceleration = dv/dt or the rate of change of v the speed was changing positively so acceleration was positive force = m*a =mass is positive and acc is positive so force is positive

OpenStudy (swissgirl):

if the backward force is greater than you would be decelerating (slowing down)

OpenStudy (swissgirl):

There is a net force on the car that is pushing the car forward making it speed up.